### v1

# DM3D: Deformable Mamba via Offset-Guided Gaussian Sequencing for Point Cloud Understanding

Abstract
State Space Models (SSMs) demonstrate significant potential for long-sequence modeling, but their reliance on input order conflicts with the irregular nature of point clouds. Existing approaches often rely on predefined serialization strategies, which cannot adjust based on diverse geometric structures. To overcome this limitation, we propose DM3D, a deformable Mamba architecture for point cloud understanding. Specifically, DM3D introduces an offset-guided Gaussian sequencing mechanism that unifies local resampling and global reordering within a deformable scan. The Gaussian-based KNN Resampling (GKR) enhances structural awareness by adaptively reorganizing neighboring points, while the Gaussian-based Differentiable Reordering (GDR) enables end-to-end optimization of serialization order. Furthermore, a Tri-Path Frequency Fusion module enhances feature complementarity and reduces aliasing. Together, these components enable structure-adaptive serialization of point clouds. Extensive experiments on benchmark datasets show that DM3D achieves state-of-the-art performance in classification, few-shot learning, and part segmentation, demonstrating that adaptive serialization effectively unlocks the potential of SSMs for point cloud understanding.
1 Introduction
Figure 1: Illustration of our deformable scanning. The offset network predicts spatial offsets $\Delta p$ and sequential offsets $\Delta t$. Guided by the predicted offsets, a Gaussian kernel performs consistent local resampling and global reordering, yielding structure-aware sequences that capture fine-grained geometric details.
3D point clouds serve as a sparse yet efficient representation of real-world geometry and are widely used in numerous vision applications, such as autonomous driving [35, 48, 46], robotics [5], and AR [18]. Unlike structured 2D images, point clouds lack a regular grid structure, meaning there is no inherent spatial adjacency between points, which poses fundamental challenges to effective representation and semantic understanding [3, 51, 17, 36, 49]. Recently, Mamba enhanced State Space Models (SSMs) [11] with a selective mechanism, achieving remarkable computational efficiency and strong long-sequence modeling capability. However, applying Mamba [10] to point clouds requires a crucial step—serialization—which plays a key role in enabling the model to capture the underlying spatial structure.
A variety of strategies have been explored to serialize point clouds for sequential modeling, including space-filling curves (SFCs) [49, 17, 43], MLP-based projections [13, 20], and voxel- or graph-based unfoldings [34, 6]. Despite their differences, these approaches share a common limitation: once serialization is completed, the token order becomes static and remains agnostic to both the input geometry and the downstream task. Such rigidity simplifies token feeding into SSMs but fundamentally contradicts the irregular and permutation-invariant nature of point clouds. Consequently, spatially adjacent points may be scattered throughout the sequence, disrupting local continuity and hindering the capture of meaningful structural details, especially in sparse or complex regions.
In contrast, human vision adaptively attends to geometric edges, local details, and structural cues, extracting the most informative patterns under limited attention. This content-aware perception efficiently forms discriminative and relational representations. Inspired by this mechanism, we inquire: can a model dynamically adjust its scanning path to the geometric structure of a point cloud, thereby endowing SSMs with genuine structural awareness? A recent effort, DefMamba [22], takes a initial step in this direction by incorporating deformable mechanisms into SSMs, where spatial and sequential offsets are learned to adjust scanning paths. However, its design is confined to 2D regular grids, where scanning corresponds merely to reordering predefined token indices [14, 24, 54, 42]. Point clouds, in contrast, lack an inherent order. Thus, deformable scanning is not merely a simple reordering of tokens, but a structure-aware sequence construction that demands joint spatial adaptivity and dynamic serialization.
To this end, we propose DM3D (Deformable Mamba for 3D point clouds), which, to our knowledge, is the first work to integrate a deformable mechanism tailored for point cloud understanding into Mamba. The core idea is to employ learnable spatial and sequential offsets to guide the Gaussian weighting, thereby achieving dual-domain deformation that unifies spatial and sequential adaptivity (see Fig. 1). Our key contributions are:
  • We introduce a deformable mechanism that dynamically adapts scanning paths based on input geometry, enhancing the capture of fine-grained geometric details.
  • We propose an offset-guided Gaussian sequencing mechanism that unifies local resampling and global reordering within the deformable scan. Its Gaussian-based KNN Resampling (GKR) enhances local structural awareness through adaptive neighbor aggregation, while the Gaussian-based Differentiable Reordering (GDR) enables end-to-end optimization of token sorting.
  • We propose a Tri-Path Frequency Fusion (TPFF) module that adaptively exchanges information across paths and suppresses information redundancy.
2 Related Work
Serialization and Mamba for Point Cloud. Point cloud serialization aims to convert unordered point sets into structured sequences compatible with sequence models. Previous methods generate sequences based on predefined traversal rules and process them with Transformer-based architectures [39, 6, 25]. As point cloud sequences become longer, Transformers [32] encounter significant computational bottlenecks. In contrast, Mamba, with its linear complexity, can handle such sequences efficiently [44, 13, 43, 27]. However, its effectiveness depends on maintaining the adjacency of semantically or spatially related points in the sequence [26, 43, 47, 16]. Thus, serialization serves as a crucial bridge between point clouds and Mamba, and different serialization schemes may be better suited to different tasks.
PointMamba [17] employs Hilbert and transposed Hilbert curves for efficient sequence generation, preserving spatial locality and providing a simple yet effective baseline. PCM [49] introduces a consistent traversal strategy, generating multiple sequence variants along coordinate axes and optimizing local continuity using order prompts. VoxelMamba [47] serializes voxelized scenes using Hilbert curves to maintain neighborhood consistency, while GridMamba [43] fuses multiple SFCs and randomly alternates them across layers to alleviate locality loss.
Despite their flexibility, these approaches rely on static, predefined mapping rules. In contrast, our DM3D treats serialization as a learnable module that allows the model to dynamically optimize point sorting during training, effectively reformulating serialization as a deformable process.
Deformable Mechanisms in Visual Tasks. Deformable mechanisms enable models to dynamically adjust their receptive fields according to input content. The concept originated from DCN [7], which introduced learnable offsets to adapt convolutional sampling locations. DAT [41] later extended this idea into the attention mechanism. In point cloud understanding, PointConv [37] incorporated deformable continuous convolution to handle non-uniform point distributions, while KPConv [30] further learned kernel displacements to better adapt to diverse local geometries. 3D-GCN [19] combined deformable kernels with graph convolution to achieve translation and scale invariance. CenterFormer [53] and PTv2 [38] respectively leveraged deformable attention for efficient multi-scale feature aggregation and spatial alignment in detection and segmentation tasks, achieving notable performance gains. Previous deformable designs have enabled CNNs and Transformers to adapt sampling positions in the spatial domain. In contrast, DM3D extends deformability to the sequence domain, enabling SSMs to adapt the serialization order based on point cloud geometry.
3 Method
3.1 Preliminary
Figure 2: Overview of DM3D. (a) Overall architecture showing the embedding, encoder, and decoder structures. (b) The Deformable Mamba Block (DMB) consists of three SSM branches: the standard forward SSM [42] (F-SSM) branch, the channel-flip backward SSM [13] (C-SSM) branch, and the deformable SSM (D-SSM) branch. (c) Deformable Scan, the core of D-SSM, predicts spatial and sequential offsets via OffsetNet, enabling unified Gaussian-based resampling and differentiable reordering. ”Spatial flow” and ”Sequential flow” indicate operations in the spatial and sequence domains, respectively. (d) LCFA provides local contextual cues for the offset network. (e) GDR performs differentiable token reordering according to learned offsets. Symbols: CatCat and $Concat$ denote concatenation along the channel dimension, $\odot$ element-wise multiplication, $\oplus$ residual addition, $\otimes$ matrix multiplication, and $\sum$ summation.
State Space Models (SSMs) are sequence modeling frameworks derived from linear time-invariant systems in control theory, capturing dependencies through latent state transitions. To adapt SSMs for discrete sequence modeling, Structured SSMs [12] (S4) employ the Zero-Order Hold (ZOH) method for discretization, introducing a timescale parameter $\Delta$. The discretized equations are:
$\bar{A}=\exp(\Delta A),\bar{B}=\Delta{{A}^{-1}}(\exp(\Delta A)-I)\Delta B$ (1)
${{h}_{t}}=\bar{A}{{h}_{t-1}}+\bar{B}{{u}_{t}},{{y}_{t}}=C{{h}_{t}}$ (2)
where $\bar{A}\in\mathbb{R}^{L\times L}$ and $\bar{B}\in\mathbb{R}^{L\times 1}$ are the discretized counterparts of continuous state evolution matrix $A$ and input projection matrices $B$, respectively. $C$ denotes the output projection matrix, and $u_{t}$, $h_{t}$, and $y_{t}$ represent the input, hidden state, and output at time step $t$.
This discrete formulation enables SSMs to model long sequences with linear complexity. Mamba [10] enhances S4 through a Selective Scan (S6) mechanism that renders $\Delta$, $B$, and $C$ input-dependent, allowing for dynamic adjustment. Coupled with the introduction of a global convolution for efficiency, these improvements establish Mamba as a robust and efficient framework for sequential processing.
3.2 Overview
DM3D adopts a variant of the ViT [15] as its backbone architecture, following the design adopted by baseline networks such as PointBERT [45] and PointMamba [42]. The overall framework is illustrated in Fig. 2(a).
Given an input point cloud, we first apply Farthest Point Sampling (FPS) and K-Nearest Neighbor (KNN) grouping to partition it into $N$ local groups, each containing $K$ points. A lightweight PointNet extracts local features $F\in{{\mathbb{R}}^{N\times D}}$, with group centers $P\in{{\mathbb{R}}^{N\times 3}}$, where $D$ denotes the embedding dimension. To inject geometric priors, the absolute coordinates of each center are encoded via an MLP to obtain position embeddings $pos\in{{\mathbb{R}}^{N\times D}}$, which are added to the local features. A classification token ${{F}_{cls}}\in{{\mathbb{R}}^{1\times D}}$ is prepended, and the resulting embeddings are serialized along a Hilbert curve to form the initial token sequence ${{X}_{0}}\in{{\mathbb{R}}^{(N+1)\times D}}$ for subsequent modeling. Each encoder stage in DM3D follows a Transformer-like architecture consisting of a Geometry-Feature Coupled Pooling (GFCP) module and a Deformable Mamba Block. The GFCP, following HyMamba [20], serves as a local enhancement layer, while the Deformable Mamba Block (DMB) performs deformable state-space modeling. The computation at the i-th stage is formulated as:
${{X}_{i}}^{\prime}=GFCP(LN({{X}_{i-1}}))+{{X}_{i-1}},$ (3)
${{X}_{i}}=DMB(LN({{X}_{i}}^{\prime}))+{{X}_{i}}^{\prime}$ (4)
where ${{X}_{i}}\in{{\mathbb{R}}^{(N+1)\times D}}$ is the output of the i-th stage and $LN$ denotes Layer Normalization.
After $T$ encoding stages, the classification token and the average-pooled feature from the final stage are concatenated and fed into a task-specific head.
3.3 Deformable Mamba Block
The Deformable Mamba Block serves as the core of this work, aiming to dynamically adjust the scanning path by reorganizing spatial and sequential structures. As illustrated in Fig. 2(b), following prior works [44, 43, 22], we replace the original 1D causal convolution with depthwise convolution and introduce a D-SSM branch, while retaining the C-SSM branch from Mamba3D [13] and the F-SSM [42] branch to maintain model stability.
Local Context Feature Aggregation (LCFA). The feature representation of a single center point is often insufficient, as geometric relationships in point clouds primarily reside within local neighborhoods. To predict meaningful offsets, it is essential to perceive the surrounding geometric context. As shown in Fig. 2(c), we aggregate features within each neighborhood to capture rich local contextual dependencies.
[…]
The center and neighborhood features are concatenated and passed through a $\times1$ convolution to produce aggregation weights (see Fig. 2(d)):
${{F}_{context}}=\sum\nolimits_{j\in\mathcal{N}(i)}{Conv_{1\times 1}([{{F}_{E}};{{F}_{\mathcal{N}(i)}}])\cdot{{F}_{\mathcal{N}(i)}}}$ (5)
where ${{F}_{E}}\in{{\mathbb{R}}^{N\times{{K}_{q}}\times D}}$ is the center feature expanded for alignment, and ${F}_{context}$} denotes the aggregated local contextual representation.
While ${F}_{context}$} encodes local geometric relationships, it lacks a global understanding of the shape. Relying solely on local features may lead to biased offset predictions. Therefore, we further concatenate it with global features to form the aggregated representation ${F}_{agg}$}, which serves as the input to the offset network.
Offset-guided Gaussian Weighting. Following the design of DAT [41], we adopt a depthwise separable convolution with a large kernel, followed by ReLU activation and a $\times1$ convolution, to predict the initial four-dimensional offsets. A channel attention (CA) module is applied after the convolution to reduce feature redundancy [22]. To prevent excessively large deformations, the offsets are bounded by a hyperbolic tangent function and then decomposed into spatial and sequential components:
${{{\mathrm{O}}}_{offset}}=Con{{v}_{1\times 1}}(ReLU(CA(DWConv({{F}_{agg}}))))$ (6)
${\mathrm{O}}{{{}^{\prime}}_{offset}}=\tanh({{{\mathrm{O}}}_{offset}})=[\Delta p;\Delta t]$ (7)
where $\Delta p\in{{\mathbb{R}}^{N\times 3}}$ represents spatial offsets for coordinate deformation, and $\Delta t\in{{\mathbb{R}}^{N\times 1}}$ denotes sequential offsets for sequence reordering.
The offset operation results in sampled positions that no longer align with the original point set, necessitating both spatial resampling and sequential reordering. We observe that these two operations inherently follow a common principle: points with higher similarity should exert stronger influence during feature aggregation and sequence organization, while the contribution of dissimilar points should gradually decay. Motivated by this, we unify resampling and reordering into a single formulation rather than treating them as independent processes. Specifically, we introduce a mechanism that applies Gaussian kernels of identical form across both spatial and sequential domains, differing only in their operational domains and neighborhood definitions. This unified formulation enables consistent weighting and smoothing for both local resampling and global reordering.
The general Gaussian kernel is defined as:
$\mathcal{W}(d;\sigma)=\exp(-\frac{{{d}^{2}}}{2\sigma^{2}})$ (8)
where $d$ is the coordinate or index distance, and $\sigma$ is the corresponding learnable scale for resampling and reordering.
Gaussian-based KNN Resampling (GKR). For resampling, after applying coordinate offsets, each offset point $p^{\prime}_{i}$ retrieves its $K_{r}$ nearest neighbors from the original point set, denoted as the index set $\mathcal{N}_{r(i)}={j_{1},\ldots,j_{K_{r}}}$. A batch mask is used to assign infinite distances between points from different batches, ensuring that neighbor search occurs only within the same batch.
For each offset point $p{{}^{\prime}}_{i}$, feature interpolation is performed within its KNN neighborhood $\mathcal{N}_{r(i)}$:
$p{{{}^{\prime}}_{i}}={{p}_{i}}+\Delta{{p}_{i}}$ (9)
$f{{}^{\prime}}_{i}^{(s)}=\sum\limits_{j\in\mathcal{N}_{r(i)}}{\frac{\mathcal{W}(\parallel{{p^{\prime}}_{i}}-{{p}_{j}}|{{|}_{2}};{{\sigma}_{s}})}{\sum\limits_{l\in\mathcal{N}_{r(i)}}{\mathcal{W}(\parallel{{p^{\prime}}_{i}}-{{p}_{l}}{{\parallel}_{2}};{{\sigma}_{s}})+\varepsilon}}}{{f}_{i}}$ (10)
where ${p}_{j}$} denotes a neighboring points, $f{{}^{\prime}}_{i}^{(s)}$ is the resampled feature, ${f}_{i}$} is the original feature at ${p}_{i}$}, ${\sigma}_{s}$} is the Gaussian scale parameter for resampling, and $\varepsilon$ is a small constant to avoid division by zero.
Notably, we deliberately omit the relative offset bias matrix in the resampling stage. In 2D image deformable convolutions, offset operations disrupt the regular grid topology, thus requiring explicit bias compensation to preserve local geometric consistency [41, 22]. However, point clouds inherently lack a fixed grid structure. Rather than imposing biased terms, we allow the network to directly learn from the raw coordinate geometry, enabling more natural spatial deformation.
Gaussian-based Differentiable Reordering (GDR). Traditional sorting operations can explicitly reorder point indices but are inherently non-differentiable, blocking gradient flow and preventing end-to-end optimization of sequence sorting [22, 9]. To overcome this limitation, we reinterpret sorting as a continuous probabilistic mapping and design the Gaussian-based differentiable reordering mechanism according to Eq. 8, which maps discrete ranking relations into a continuous probabilistic space. Instead of assigning each token to a single discrete index, its sorting position is represented as a continuous Gaussian-weighted distribution over adjacent sequence locations, as illustrated in Fig. 2(e).
Specifically, we add the token offsets $\Delta t$ to the base index vector $I\in{{\mathbb{R}}^{N\times 1}}$ (derived from the Hilbert curve) to obtain the shifted index vector $s$. We then compute the pairwise difference between $s$ and the target index vector $J=[1,2,\ldots,N]$, forming a difference matrix $(s-J)\in{{\mathbb{R}}^{N\times N}}$. This matrix encodes the relative positional deviation between tokens, which is then used to perform Gaussian-weighted aggregation for differentiable reordering:
${{s}_{i}}={{I}_{i}}+\Delta{{t}_{i}},$ (11)
$f{{}^{\prime}}_{i}^{(t)}=\sum\limits_{j=1}^{N}{\frac{\mathcal{W}({{s}_{i}}-{{J}_{j}};{{\sigma}_{t}})}{\sum\nolimits_{l=1}^{N}{\mathcal{W}({{s}_{i}}-{{J}_{l}};{{\sigma}_{t}})+\varepsilon}}}F{{}^{\prime}}_{i}^{(s)}$ (12)
where $f{{}^{\prime}}_{i}^{(t)}$ denotes the reordered $i$-th token feature, $F{{}^{\prime}}^{(s)}$ is the global resampled feature matrix obtained by stacking all local resampled features $\{f_{i}^{\prime(s)}\}_{i=1}^{N}$ from the GKR, and $\sigma_{t}$ is the learnable Gaussian scale parameter controlling the smoothness of sorting.
During Gaussian-weighted aggregation, token features are softly distributed across neighboring sequence positions, forming a continuous and differentiable relaxation of discrete sorting. Importantly, GDR does not replace non-differentiable sorting but fully subsumes it within a continuous framework (See Appendix for detailed derivation.):
  • When ${\sigma}_{t}$$\to$$\infty$, GDR degenerates into global average pooling, with gradients vanishing as the derivative converges to zero everywhere.
  • For any fixed ${{\sigma}_{t}}>0$, the normalized Gaussian weights remain continuous and differentiable with respect to ${s}_{i}$} , ensuring smooth gradient propagation and enabling end-to-end optimization.
  • As ${{\sigma}_{t}}\to{{0}^{+}}$, the mapping converges to deterministic sorting, where each position ${s}_{i}$} inherits the feature of its nearest discrete index ${J}_{j}$}, effectively snapping tokens to their closest sequence positions.
Although token indices are not explicitly permuted, their features are nearly exchanged, achieving an effect equivalent to differentiable sorting. By controlling $\sigma_{t}$, GDR provides a smooth approximation to discrete ranking—smaller $\sigma_{t}$ approximates standard sorting, while larger values approach global averaging—thus preserving the end-to-end learnability of token reordering.
In summary, deformable scanning enhances point cloud understanding by learning offsets that adaptively adjust point coordinates and scanning order across spatial and sequential domains. This design endows the model with strong geometric adaptability.
3.4 Tri-Path Frequency Fusion
In the DMB, the F-SSM, C-SSM, and D-SSM capture dynamic features along different pathways. Directly fusing these paths in the spatial domain may cause information redundancy, as signals of different paths overlap. To address this, we map the sequential features into the frequency domain, where cross-path information can be more effectively complemented (see Fig. 3).
Figure 3: Illustration of TPFF. Cross-path fusion is performed first, followed by frequency enhancement.
Table 1: Classification on ModelNet40 and ScanObjectNN. We report overall accuracy (%), number of parameters (#P), and FLOPs (#F). $\dagger$ denotes voting strategy.
Reference Methods
ModelNet40
1k P
ScanObjectNN #P(M) #F(G)
OBJ_BG OBJ_ONLY PB_T50_RS
Supervised Learning Only
CVPR 17 PointNet [3] 89.2 73.3 79.2 68.0 3.5 0.5
NeurIPS 17 PointNet++ [29] 90.7 82.3 84.3 77.9 1.5 1.7
TOG 19 DGCNN [36] 92.9 82.8 86.2 78.1 1.8 2.4
JAS 23 PointConT [23] 93.5 - - 90.30 - -
NeurIPS 24 PointMamba [17] - 88.30 87.78 82.48 12.3 3.6
AAAI 25 PCM [49] 93.4 - - 88.1 34.2 45.0
CVPR 25 SAST [1] 92.7 92.25 91.39 87.30 12.3 3.6
DM3D 93.76 90.70 91.56 89.40 13.6 4.4
With Self-supervised Pre-training
CVPR 22 Point-BERT [45] 93.4 88.12 88.30 83.69 22.1 4.8
ECCV 22 Point-MAE [28] 94.4 91.22 90.02 84.94 22.1 4.8
NeurIPS 23 PointGPT-S [4] 94.0 91.6 90.0 86.9 29.2 5.7
CVPR 24 PointDif [52] - 93.29 91.91 87.61 - -
NeurIPS 24 PointMamba [17] 93.6 94.32 92.60 89.31 12.3 3.6
ACM MM 24 Mamba3D [13] 95.1 95.18 92.60 93.34 16.9 3.9
NeuCom 25 E-Mamba [16] 94.7 94.32 92.94 91.98 13.78 7.49
ICCV 25 Point-PQAE [50] 94.0 95.0 93.6 89.6 22.1 -
CVPR 25 SAST [1] 93.4 94.32 91.91 89.10 12.3 3.6
DM3D 94.53 94.32 93.72 93.09 13.6 4.4
DM3D 95.11 95.21 93.93 94.32 13.6 4.4
The three pathway features, $F_{F},F_{C},F_{D}\in\mathbb{R}^{N\times D}$, are first adaptively modulated by the other two branches before cross-fusion:
$w_{i}=\text{sigmoid}(F_{i}),\quad F^{\prime}_{i}=F_{i}\odot\frac{1}{2}(w_{j}+w_{k})$ (13)
where $(i,j,k)$ cyclically traverses $\{(F,C,D),$ $(C,D,F),$ (D,F,C)}(D,F,C)\}. This enables each pathway to perceive the activation strength of the others, allowing dynamic adjustment of their relative importance.
The modulated features are concatenated along the channel dimension (yielding $3D$ channels) and fused via a grouped convolution to reduce dimensionality back to $D$. To maintain inter-group communication, channel shuffle is applied after fusion:
$F_{shuffle}=\text{shuffle}(Conv_{1\times 1}([F^{\prime}_{F};F^{\prime}_{C};F^{\prime}_{D}]))$ (14)
We further perform adaptive enhancement in the frequency domain. Specifically, a Fast Fourier Transform (FFT) is applied along the sequence dimension, followed by a $\times1$ convolution on the real and imaginary components for frequency modulation, and finally, an inverse FFT (IFFT) reconstructs the spatial features. The overall process is formulated as:
$F_{FFT}=\mathcal{F}(F_{shuffle})=\Re(F_{FFT})+i\Im(F_{FFT})$ (15)
$[F_{FFT}^{r};F_{FFT}^{i}]=Conv_{1\times 1}([\Re(F_{FFT});\Im(F_{FFT})])$ (16)
$F_{out}=\Re\left(\mathcal{F}^{-1}(F_{FFT}^{r}+iF_{FFT}^{i})\right)$ (17)
where $\mathcal{F}$ and ${\mathcal{F}}^{-1}$} denote the FFT and IFFT, respectively. $\Re$ and $\Im$ represent the real and imaginary parts.
The TPFF projects spatial features from all SSM branches into a unified frequency domain to mitigate aliasing across overlapping receptive fields.
4 Experiments
4.1 Implementation Details
All experiments are conducted on a single NVIDIA TITAN RTX GPU. We set the total number of stages $T$ to 6, the hidden dimension $D$ to 384, the number of points $N$ sampled by FPS to 128, and the number of neighbors $K$ in KNN to 32. In the GKR, $K_{r}$ is set to 3 and $\sigma_{s}$ to 1, while in the GDR, $\sigma_{t}$ is set to 0.2. The models are trained for 300 epochs using the AdamW optimizer with the cosine learning rate scheduler (CosLR), an initial learning rate of 5e-4, weight decay of 0.05, warmup for 10 epochs, and batch size of 32. More parameter details are provided in the Appendix. The pre-training is conducted on ShapeNetCore [2], following prior protocols [28, 45, 42, 13]. Both classification and part segmentation adopt the cross-entropy loss.
4.2 Downstream Tasks
Classification. We adopt a three-layer MLP as the classification head and evaluate our method on ModelNet40 [40] and ScanObjectNN [31]. ModelNet40 uses 1024 input points with scale&translate augmentation, while ScanObjectNN uses 2048 points with rotation augmentation. As shown in Tab. 1, DM3D achieves highly competitive results. Without pre-training, it achieves 93.76% on ModelNet40 and 89.40% on PB_T50_RS, outperforming baselines such as PointMamba [17] and PCM [49], demonstrating strong robustness to perturbations. With pre-training, DM3D attains 93.09% overall accuracy (OA) on PB_T50_RS without voting, surpassing recent counterparts, including Point-PQAE [50] (89.6%) and SAST [1] (89.1%). Moreover, DM3D remains significantly more lightweight than PointGPT-S [4] and PCM [49], with only a minimal increase in computational cost.
Few-shot Learning. We conduct few-shot classification experiments following prior work [17, 13, 45, 20] to further validate the few-shot learning capability of DM3D. As shown in Tab. 2, DM3D consistently outperforms other methods under the purely supervised setting, demonstrating strong feature learning capability. With pre-training, DM3D remains highly competitive, matching or surpassing Transformer-based methods, which highlights its robustness and generalization.
[…]
Methods 5-Way 10-Way
10-Shot 20-Shot 10-Shot 20-Shot
Supervised Learning Only
PointNet [3] 52.03.8 57.84.9 46.64.3 35.24.8
DGCNN [36] 31.62.8 40.84.6 19.92.1 16.91.5
OcCo [33] 90.62.8 92.51.9 82.91.3 86.52.2
DM3D 91.06.9 94.83.6 86.15.5 91.753.2
With Self-supervised Pre-training
ACT [8] 96.82.3 98.01.4 93.34.0 95.62.8
Point-BERT [45] 94.63.1 96.32.7 91.05.4 92.75.1
Point-MAE [28] 96.32.5 97.81.8 92.64.1 95.03.0
PointGPT-S [4] 96.82.0 98.61.1 92.64.6 95.23.4
PointMamba [17] 95.02.3 97.31.8 91.44.4 92.84.0
DM3D 95.94.5 97.5±3.2 91.255.1 95.73.0
Table 3: Part segmentation on the ShapeNetPart. We reports class-level mIoU (Cls.mIoU) and instance-level mIoU (Ins.mIoU).
Method Cls.mIoU(%) Ins.mIoU(%)
Supervised Learning Only
PointNet [3] 80.4 83.7
PointNet++ [29] 81.9 85.1
DGCNN [36] 82.3 85.2
DM3D 83.7 85.3
With Self-supervised Pre-training
OcCo [33] 83.4 84.7
ACT [8] 84.7 86.2
MaskPoint [21] 84.6 86.0
Point-BERT [45] 84.1 85.6
Point-MAE [28] 84.2 86.1
PointMamba [17] 84.4 86.0
Mamba3D [13] 83.6 85.6
Point-PQAE [50] 84.6 86.1
DM3D 84.6 86.2
Part Segmentation. ShapeNetPart [2] contains 50 part categories across 16 object classes. We adopt 2048 input points without normals, and the segmentation head follows a PointNet++-like [29] design that extracts features from the 2nd, 4th, and 6th encoder layers. Tab. 3 shows that ShapeNetPart is extremely challenging, and even state-of-the-art methods only marginally improve compared to each other. With pre-training, DM3D achieves 84.6% Cls.mIoU and 86.2% Ins.mIoU. This surpasses the PointMamba baseline by 0.2% and matches or slightly outperforms recent methods such as Mamba3D [13] and Point-PQAE [50], validating that our deformable mechanism strengthens SSMs in capturing fine-grained geometric details. Visual results are provided in the Appendix.
4.3 Ablation Studies
To comprehensively evaluate the effectiveness of each component, we conduct ablation studies on the model architecture and its variants, with all trained from scratch.
Effect of Each Component. As shown in Tab. 4, the full model significantly outperforms all ablated variants on both tasks. Removing the deformable branch (w/o D-SSM) or the frequency fusion module (w/o TPFF) leads to performance drops of 3.28%/3.95% and 0.82%/0.88%, respectively, confirming the effectiveness of the deformable mechanism and the importance of cross-path feature interaction.
We further ablate the components of deformable scanning. Removing LCFA causes performance drops of 1.3% and 0.47% on OBJ_ONLY and PB_T50_RS respectively, indicating that neighborhood aggregation provides richer contextual cues for offset prediction. Eliminating coordinate offsets causes a 2.85% drop, showing that spatial deformation enhances structural perception. Removing the GKR leads to 2.4% degradation, as the model loses the ability to capture new local structures. Similarly, removing the GDR yields a 1.4% performance loss, as it fails to prioritize scanning similar structures. These results demonstrate that our method improves performance in point cloud understanding, validating our design choices.
Table 4: Ablation on model architecture and components of deformable scanning.
Method OBJ_ONLY PB_T50_RS #P (M) #F (G)
(a) Ablation on model architecture
$\star$ Full 91.56 89.40 13.6 4.4
w/o D-SSM 88.28 85.45 8.89 3.04
w/o TPFF 90.74 88.52 13.39 4.38
(b) Ablation on deformable scanning components
w/o LCFA 90.26 88.93 12.71 3.47
w/o OffsetNet 88.71 86.93 13.59 4.4
w/o GKR 89.12 86.08 13.6 4.4
w/o GDR 89.93 88.12 13.59 4.4
Table 5: Ablation on key parameters and fusion strategies. We evaluate the effects of neighborhood size $K_{r}$ in GKR, reordering methods and Gaussian scale $\sigma_{t}$ in GDR, and path fusion strategies.
Variants OBJ_ONLY PB_T50_RS
(a) Neighborhood size $K_{r}$ in GKR.
$K_{r}=2$ 91.40 89.27
$K_{r}=3$ (ours) 91.56 89.40
$K_{r}=4$ 91.35 89.21
$K_{r}=5$ 91.23 89.33
(b) Reordering methods and Gaussian scale $\sigma_{t}$ in GDR.
Fixed order (Hilbert) 89.93 88.12
Non-differentiable sorting 90.81 89.01
GDR ($\sigma_{t}{=}0.1$) 91.35 89.32
GDR ($\sigma_{t}{=}0.2$, ours) 91.56 89.40
GDR ($\sigma_{t}{=}0.4$) 91.23 89.29
GDR ($\sigma_{t}{=}0.6$) 91.12 89.04
(c) Path fusion strategies.
Element-wise mean 90.74 88.52
Learned linear fusion 90.05 88.63
Convolutional fusion 89.86 87.42
TPFF (ours) 91.56 89.40
Effect of resampling neighborhood size $K_{r}$. $K_{r}$ defines the resampling neighborhood after offset. Smaller values cause insufficient local aggregation, while larger ones introduce irrelevant points. As shown in Tab. 5(a), $K_{r}=3$ offers the best trade-off. In contrast, $K_{q}$ in LCFA only indirectly influences final performance, and the computational cost increases linearly with its value. Therefore, $K_{q}=8$ is fixed for efficiency.
Effect of reordering methods and scale ${\sigma}_{t}$. Removing reordering or replacing GDR with a non-differentiable sorting operation causes 1.63%/1.28% and 0.75%/0.36% drops, respectively, confirming the necessity of dynamic sequence adjustment. While non-differentiable sorting introduces index reordering, its lack of gradient flow makes it inferior to GDR, which directly optimizes token shifts. Among GDR, the best result is achieved at ${{\sigma}_{t}}=0.2$, where smaller values make the sorting overly strict and larger ones oversmooth discriminative features.
Effect of fusion schemes. We compare multiple path fusion strategies and observe that frequency fusion consistently surpasses spatial ones. Element-wise mean and linear fusions offer limited capacity for cross-path complementarity, while convolutional fusion insufficiently mitigates aliasing.
4.4 Visualization and Analysis
To better demonstrate the working effect of the deformable scan, we perform a visual analysis as shown in Fig. 4.
(a) Point coordinates before and after deformation.
(b) Token feature interaction matrix (log-scale) of GDR.
(c) Feature contribution sources for tokens at indices 40 and 80.
Figure 4: Visualization of the deformable mechanism. The token feature interaction matrix shows the feature weights from source tokens (x-axis) to target tokens (y-axis).
Analysis of Coordinate Offset. Larger offsets are observed near object boundaries, where richer geometric variations exist (see Fig. 4(a)). This indicates that the model learns to shift points toward—and focus resampling on—these semantically salient areas, thereby enhancing its sensitivity to fine structural details.
Analysis of Differentiable Reordering. The heatmap in Fig. 4(b) shows that each token is assigned a continuous Gaussian-weighted distribution over target positions rather than a single discrete index, confirming that GDR performs probabilistic rather than deterministic reordering. Most weights concentrate along the main diagonal, indicating that GDR preserves global stability while allowing local flexibility; scattered off-diagonal activations indicate feature correlations between points. Target points receive contributions from neighboring points that share geometric or semantic similarities (see Fig. 4(c)). This encourages similar points to occupy adjacent sequence positions and prevents structural disruption caused by sorting, thereby maintaining local consistency.
5 Conclusion
In this paper, we propose DM3D, a deformable Mamba network for point cloud understanding. Unlike previous models that rely on fixed serialization, DM3D introduces an offset-guided deformable mechanism with Gaussian weighting, unifying local resampling and global reordering, and enabling end-to-end optimization of geometric structures and scanning paths. Extensive experiments on ModelNet40, ScanObjectNN, and ShapeNetPart demonstrate state-of-the-art performance, confirming that point cloud serialization should be learned and adapted as part of the model rather than treated as a static preprocessing step. While DM3D achieves compelling results, it introduces moderate efficiency overhead on extremely large-scale point clouds owing to the deformable operations. Future work will explore more lightweight formulations to improve efficiency in large-scale scenarios.

### v2

@@
 Extensive experiments on benchmark datasets show that DM3D achieves state-of-the-art performance in classification, few-shot learning, and part segmentation, demonstrating that adaptive serialization effectively unlocks the potential of SSMs for point cloud understanding.
+The code will be released at https://github.com/L1277471578/DM3D
 1 Introduction
@@
 Future work will explore more lightweight formulations to improve efficiency in large-scale scenarios.
+Supplementary Material
+6 Analysis of GDR Differentiability
+Figure 5: Visual results of part segmentation by DM3D and PointMamba on ShapeNetPart.
+In this section, we analyze the Gaussian weights and their derivatives with respect to the offset indices $s_{i}$ to examine the behavior of Gaussian-based Differentiable Reordering (GDR) mechanism.
+The weighting function is defined as:
+$W_{ij}^{(t)}=\frac{\exp\left(-\frac{{{({{s}_{i}}-{{J}_{j}})}^{2}}}{2\sigma_{t}^{2}}\right)}{\sum\nolimits_{l=1}^{N}{\exp\left(-\frac{{{({{s}_{i}}-{{J}_{l}})}^{2}}}{2\sigma_{t}^{2}}\right)}}$ (18)
+where ${\sigma}_{t}$} is the Gaussian scale in the sequence domain, $J=[1,2,\ldots,N]$ is the target index vector, and ${{s}_{i}}={{I}_{i}}+\Delta{{t}_{i}}$ denotes the index of the $i$-th token after applying the learned offset $\Delta{{t}_{i}}$.
+To analyze gradient flow, the first-order partial derivative of $W_{ij}^{(t)}$ with respect to $s_{i}$ is given by:
+$\frac{\partial W_{ij}^{(t)}}{\partial{{s}_{i}}}=\frac{W_{ij}^{(t)}}{{{\sigma}_{t}}^{2}}\left({{J}_{j}}-\sum\limits_{l=1}^{N}{W_{il}^{(t)}}{{J}_{l}}\right)$ (19)
+We analyze three distinct cases: $\sigma_{t}\to+\infty$, $\sigma_{t}>0$, and $\sigma_{t}\to 0^{+}$
+Case1: Analysis of ${\sigma}_{t}$$\to$$\infty$. As $2\sigma_{t}^{2}\to+\infty$, the exponential terms converge to 1, yielding:
+$\underset{{{\sigma}_{t}}\to+\infty}{\mathop{\lim}}\,W_{ij}^{(t)}=\frac{1}{N}$ (20)
+$\underset{{{\sigma}_{t}}\to+\infty}{\mathop{\lim}}\,\frac{\partial W_{ij}^{(t)}}{\partial{{s}_{i}}}=0$ (21)
+In this case, GDR degenerates to global average pooling, and all gradients vanish.
+Case2: Analysis of ${{\sigma}_{t}}>0$.
+When ${\sigma}_{t}$} is finite and positive, $W_{ij}^{(t)}$ is infinitely differentiable with respect to $s_{i}$.
+Since $W_{ij}^{(t)}\in(0,1)$, its derivative remains bounded and continuous with respect to $s_{i}$.
+In this case, GDR defines a smooth and continuously differentiable mapping that supports stable gradient backpropagation through $s_{i}$.
+Case3: Analysis of ${{\sigma}_{t}}\to{{0}^{+}}$.
+According to Eq.
+18, the relationship between $s_{i}$ and the indices determines the weight, yielding two scenarios: $s_{i}$ has a single nearest neighbor, or is equidistant to two indices.
+Let $d_{\min}^{2}=\min_{l}(s_{i}-J_{l})^{2}$, and define the index set $\mathcal{T}=\{j|(s_{i}-J_{j})^{2}=d_{\min}^{2}\}$ with cardinality $m=|\mathcal{T}|$. The Gaussian kernel can be rewritten as
+$\mathcal{W}=\exp\left(-\frac{{{({{s}_{i}}-{{J}_{j}})}^{2}}}{2\sigma_{t}^{2}}\right)$ (22)
+$=\exp\left(-\frac{{{d}_{\min}}^{2}}{2\sigma_{t}^{2}}\right)\cdot\exp\left(-\frac{{{({{s}_{i}}-{{J}_{j}})}^{2}}-{{d}_{\min}}^{2}}{2\sigma_{t}^{2}}\right)$
+For all $j\in\mathcal{T}$, the second term equals 1; for $j\notin\mathcal{T}$, it decays exponentially. Substituting into the normalized weight:
+$W_{ij}^{(t)}=\frac{\exp\left(-\frac{{{({{s}_{i}}-{{J}_{j}})}^{2}}-{{d}_{\min}}^{2}}{2\sigma_{t}^{2}}\right)}{\sum\nolimits_{l\in\mathcal{T}}{1}+\sum\nolimits_{l\notin\mathcal{T}}{\exp\left(-\frac{{{({{s}_{i}}-{{J}_{j}})}^{2}}-{{d}_{\min}}^{2}}{2\sigma_{t}^{2}}\right)}}$ (23)
+Case3a: $s_{i}$ has a single nearest neighbor $J_{k}$.
+When $s_{i}$ has a single nearest neighbor $J_{k}$—either coinciding with or lying between indices—we have $\mathcal{T}={k}$, $m=1$, and $d_{\min}\geq 0$.
+The non-minimum terms vanish exponentially, yielding
+[…]
+For the derivative, both ${{W}^{(t)}}_{il}$ and its weighted sum decay exponentially, leading to
+$\underset{{{\sigma}_{t}}\to{{0}^{+}}}{\mathop{\lim}}\,\frac{\partial W_{ij}^{(t)}}{\partial{{s}_{i}}}=0,\forall J$ (25)
+In this scenario, the weight deterministically assigns $s_{i}$ to its nearest index $J_{k}$, and the gradient vanishes, indicating that GDR degenerates into exact sorting.
+Case 3b: $s_{i}$ is equidistant to two indices.
+In this scenario, $m=2$, and all $j\in\mathcal{T}$ have $d_{\min}$, while other indices have greater distances.
+As ${{\sigma}_{t}}\to{{0}^{+}}$, the minor terms in the denominator vanish, leading to:
+[…]
+For the derivative, when $j\notin\mathcal{T}$, ${{W}^{(t)}}_{ij}$ decays exponentially, so the derivative tends to zero. For $j\in\mathcal{T}$, the derivative is given by:
+$\frac{\partial{{W}^{(t)}}_{ij}}{\partial{{s}_{i}}}=\frac{1}{2{{\sigma}_{t}}^{2}}({{J}_{j}}-{{\bar{J}}_{\mathcal{T}}}),\bar{J}_{\mathcal{T}}=\frac{1}{m}\sum_{l\in\mathcal{T}}J_{l}$ (27)
+where $\bar{J}_{\mathcal{T}}$ denotes the mean index of $\mathcal{T}$, and ${{J}_{j}}-{{\bar{J}}_{\mathcal{T}}}\neq 0$. Therefore:
+$\underset{{\sigma_{t}}\to 0^{+}}{\lim}\frac{\partial W^{(t)}_{ij}}{\partial s_{i}}=\begin{cases}0,&j\notin\mathcal{T}\\ +\infty,&j\in\mathcal{T},\ J_{j}>\bar{J}_{\mathcal{T}}\\ -\infty,&j\in\mathcal{T},\ J_{j}<\bar{J}_{\mathcal{T}}\end{cases}$ (28)
+In this scenario, weights are evenly distributed the across equidistant indices, and the derivatives diverge.
+However, even a slight deviation in $s_{i}$ breaks the symmetry, restoring finite gradients (as in Case 3a).
+Consequently, the model naturally avoids such equidistant states during training.
+Overall, GDR maps each offset index $s_{i}$ to a probabilistic assignment over target indices, transforming non-differentiable sorting into a continuous, differentiable process and preserving gradient flow.
+The weighting function is differentiable for any finite ${{\sigma}_{t}}>0$, where ${\sigma}_{t}$} controls the smoothness of reordering.
+As ${{\sigma}_{t}}\to{{0}^{+}}$, it converges to non-differentiable sorting, yielding vanishing gradients in the non-equidistant scenario.
+Conversely, as $\sigma_{t}\to\infty$, the GDR degenerates into average pooling and loses its reordering capability.
+Hence, ${\sigma}_{t}$} is kept small in practice, and a minor perturbation term is introduced for stability.
+7 More Experimental Details
+Implementation Details.
+Tab.
+6 details the training and model parameters.
+We use the official pre-trained Point-MAE model.
+To avoid excessive offset from high layer stages, we set the number of stage layers to 6, stable convergence across $\sigma_{t}$ initializations (0.05–1), no gradient collapse..
+We evaluate our method on ModelNet40, ScanObjectNN, and ShapeNetPart[17].
+Table 6: Implementation details and hyperparameter settings across different datasets.
+Setting ModelNet40 ScanObjectNN ShapeNetPart
+Optimizer AdamW AdamW AdamW
+Learning rate 5e-4 5e-4 2e-4
+Weight decay 0.05 0.05 0.05
+Scheduler CosLR CosLR CosLR
+Warmup epochs 10 10 10
+Drop path rate 0.1 0.1 0.3
+Training epochs 300 300 300
+Batch size 32 32 32
+Input points 1024 2048 2048
+Hidden dim ($D$) 384 384 384
+Stage layers ($T$) 6 6 6
+FPS samples ($N$) 128 128 128
+KNN neighbors ($K$) 32 32 16
+DWConv kernel size 5 5 5
+SFC Hilbert Hilbert Hilbert
+LCFA $K_{q}$ 8 8 8
+GKR $K_{r}$ 3 3 3
+GDR $\sigma_{t}$ 0.2 0.2 0.2
+Augmentation Scale&Trans Rotation
+Part segmentation visualization.
+Fig.
+5 presents qualitative part segmentation comparisons on the ShapeNetPart validation set, including Ground Truth, the PointMamba, and our DM3D predictions.
+DM3D achieves highly competitive results, showing notable improvement on the “motorbike” category.
+8 Model details in PyTorch style pseudo-code
+We provide PyTorch-style pseudocode for the proposed modules, including the Deformable Scan for Point Clouds, Tri-Path Frequency Fusion, and Deformable Mamba Block.
+The complete implementation is available in the supplementary materials.
+Algorithm 1: Pseudo-code of the Deformable Scan For Point Cloud.
+# # Deformable Scan For Point Cloud
+class DeformableScan3DForPointCloud($n$.Module):
+def __init__(self, in_channels: int, num_neighbors: int = 8, dp_scale = 1.0, use_ca = True, resample_k = 3, enable_dp = True, enable_dt = True, expand = 2):
+super().__init__()
+self.in_channels = in_channels
+self.num_neighbors = num_neighbors
+self.dp_scale = dp_scale
+self.resample_k = resample_k
+self.temperature = temperature
+self.expand = expand
+self.enable_dp = enable_dp
+self.enable_dt = enable_dt
+self.use_ca = use_ca
+# convolution for feature aggregation
+self.dynamic_conv = $n$.Sequential($n$.Conv1d(in_channels * 2, in_channels, kernel_size=1), $n$.GELU())
+# Offset prediction network
+self.offset_net = $n$.Sequential($n$.Conv1d(in_channels * 2, in_channels, kernel_size=5, padding=2, groups=in_channels),
+$n$.Conv1d(in_channels, in_channels // 2, kernel_size=1),
+ChannelAttention(in_channels // 2) if use_ca else $n$.Identity(),
+$n$.ReLU(),
+$n$.Conv1d(in_channels//2, 4, kernel_size=1)) # 4 channels for (dx, dy, dz, dt)
+self.sigma = $n$.Parameter(torch.tensor(0.2))
+def knn_resample(self, new_coord, original_coord, original_feat, batch, k=3,sigma=0.2):
+batch_center = batch[:new_coord.size(0)] # [M]
+dists = torch.cdist(new_coord, original_coord) # [M, N]
+mask = batch_center.unsqueeze(1) != batch.unsqueeze(0)
+dists = dists.masked_fill(mask, float(’inf’))
+# Find neighbors
+topk = torch.topk(dists, k=k, dim=1, largest=False)
+weights = F.softmax(-0.5 * (topk.values / self.sigma) ** 2) # [M, k]
+knn_indices = topk.indices # [M, k]
+knn_feat = original_feat[knn_indices.view(-1)] # [M*k, C]
+knn_feat = knn_feat.view(knn_indices.size(0), knn_indices.size(1), -1) # [M, k, C]
+return (knn_feat * weights.unsqueeze(-1)).sum(dim=1) # [M, C]
+def sorting_global(features, delta_t, base_idx=None, sigma=0.2, eps=1e-8):
+B, N, C = features.shape
+device = features.device
+dtype = features.dtype
+i = base_idx.view(B, N, 1).to(dtype)
+j = torch.arange(N, device=device, dtype=dtype).view(1, 1, N).expand(B, 1, N)
+new_order = i + delta_t.view(B, N, 1)
+diff = new_order - j # [B, N, N]
+W = torch.exp(-0.5 * (diff / (sigma_t + 1e-12)) ** 2)
+W_sum = W.sum(dim=-1, keepdim=True).clamp_min(eps)
+W = W / W_sum
+permuted = torch.bmm(W, features)
+return permuted, new_order
+def forward(self, coords, feats, base_idx = None) :
+input: coords[B, N, 3], feats[B, N+1, C], base_idx[B, N]
+output: resampled_feat[B, N, C], new_coord[B, N, 3], new_order[B, N, 1]
+B, _, C = feats.shape
+device = coords.device
+cls_token = feats[:, 0, :].view(B, 1, C)
+feats = feats[:, 1:, :] # Remove class token: [B, N, C]
+N = feats.size(1)
+flat_coords = coords.view(-1, 3) # [B*N, 3]
+# LCFA
+k = min(self.num_neighbors, N - 1)
+idx = pointnet2_utils.ball_query(0.1, k, coords, coords).view(B, N, k)
+idx = idx.to(device)
+feats_trans = feats.permute(0, 2, 1).contiguous()
+group_feat = pointnet2_utils.grouping_operation(feats_trans, idx)
+group_feat = group_feat.permute(0, 2, 1, 3) # [B, N, C, k]
+center_feat = feats.unsqueeze(-1).expand(-1, -1, -1, k) # [B, N, C, k]
+local_feat = torch.cat([center_feat, group_feat], dim=2) # [B, N, 2C, k]
+local_feat_flat = local_feat.view(B * N, -1, k)
+dynamic_weights = self.dynamic_conv(local_feat_flat) # [B*N, C, k]
+context_feat = torch.sum(group_feat.contiguous().view(B * N, C, k) * dynamic_weights, dim=2) # [B*N, C]
+# Offset prediction network
+offset_input = torch.cat([feats.reshape(B * N, C), context_feat], dim=1)
+offset_input = offset_input.unsqueeze(0).permute(0, 2, 1) # [1, 2C, B*N]
+offset = self.offset_net(offset_input) # [1, 4, B*N]
+offset = offset.squeeze(0).permute(1, 0) # [B*N, 4]
+offset = torch.tanh(offset) * self.dp_scale # Normalize offsets
+delta_p = offset[:, :3] if self.enable_dp else torch.zeros_like(flat_coords)
+delta_t = offset[:, 3] if self.enable_dt else torch.zeros(B * N, device=device)
+# GKR
+new_coord = flat_coords + delta_p # [B*N, 3]
+batch_idx = torch.arange(B, device=device).repeat_interleave(N)
+resampled_feat = self.knn_resample( new_coord, flat_coords, feats.reshape(B * N, C), batch_idx, k=self.resample_k, sigma=self.sigma) # [B*N, C]
+# GDR
+resampled_feat += feats.reshape(B * N, C) # Residual connection [B*N, C]
+if self.enable_dt: # GDR
+delta_t_batched = delta_t.view(B, N) # [B, N]
+resampled_feat = resampled_feat.reshape(B, N, C) # [B, N, C]
+resampled_feat, new_order = sorting_global( resampled_feat, delta_t_batched, base_idx=base_idx, sigma=self.sigma)
+resampled_feat = torch.cat((cls_token, resampled_feat), dim=1) # [B, N+1, C]
+return resampled_feat, new_coord, new_order
+Algorithm 2: Pseudo-code of the Tri-Path Frequency Fusion.
+# # Tri-Path Frequency Fusion
+class TPFF($n$.Module):
+def __init__(self, dim: int, bias: bool = False, groups=32) -> None:
+super().__init__()
+self.fuse_proj = $n$.Conv1d(dim * 3, dim, kernel_size=1, bias=bias, groups=groups)
+self.freq_enhance = FrequencyEnhanceBlock(in_channels=dim, groups=groups)
+self.ChannelShuffle = ChannelShuffle(groups)
+def forward(self, x: torch.Tensor, y: torch.Tensor, z: torch.Tensor) -> torch.Tensor:
+input: x, y, z: F-SSM,C-SSM,D-SSM feat [B, D, L]
+output: out: fusion feat [B, D, L]
+w_x = torch.sigmoid(x)
+w_y = torch.sigmoid(y)
+w_z = torch.sigmoid(z)
+x_att = x * (w_y + w_z) / 2
+y_att = y * (w_x + w_z) / 2
+z_att = z * (w_x + w_y) / 2
+fused = torch.cat([x_att, y_att, z_att], dim=1)
+fused = self.fuse_proj(fused)
+fused = self.ChannelShuffle(fused)
+out = self.freq_enhance(fused)
+return out
+class ChannelShuffle($n$.Module):
+def __init__(self, groups):
+super().__init__()
+self.groups = groups
+def forward(self, x):
+batch, channels, length = x.size()
+channels_per_group = channels // self.groups
+x = x.view(batch, self.groups, channels_per_group, length)
+x = torch.transpose(x, 1, 2).contiguous()
+return x.view(batch, channels, length)
+class FrequencyEnhanceBlock($n$.Module):
+def __init__(self, in_channels: int, groups=32) -> None:
+super().__init__()
+self.conv = $n$.Conv1d(2 * in_channels, 2 * in_channels, kernel_size=1, groups=groups)
+def forward(self, x):
+fft_result = torch.fft.fft(x, dim=-1)
+real, imag = fft_result.real, fft_result.imag
+freq_cat = torch.cat([real, imag], dim=1)
+freq_out = self.conv(freq_cat)
+c = freq_out.shape[1] // 2
+real_new = freq_out[:, :c, :]
+imag_new = freq_out[:, c:, :]
+complex_tensor = torch.complex(real_new, imag_new)
+x_out = torch.fft.ifft(complex_tensor, dim=-1)
+return x_out.real
+Algorithm 3: Pseudo-code of the Deformable Mamba Block.
+# # Deformable Mamba Block
+class Mamba($n$.Module):
+def __init__(self, d_model, d_state=16, d_conv=4, expand=2, dt_rank="auto", dt_min=0.001, dt_max=0.1, dt_init="random", dt_scale=1.0, dt_init_floor=1e-4, conv_bias=True, use_fast_path=True, layer_idx=None, device=None, dtype=None, mamba_type="v3", enable_dp=True, enable_dt=True, num_neighbors=8): factory_kwargs = {"device": device, "dtype": dtype}
+super().__init__()
+self.d_model = d_model
+self.d_state = d_state
+self.d_conv = d_conv
+self.expand = expand
+self.d_inner = int(self.expand * self.d_model)
+self.dt_rank = math.ceil(self.d_model / 16) if dt_rank == "auto" else dt_rank
+self.use_fast_path = use_fast_path
+self.layer_idx = layer_idx
+self.mamba_type = mamba_type
+self.enable_dt = enable_dt
+self.num_neighbors = num_neighbors
+#projection layer
+self.in_proj = $n$.Linear(self.d_model, self.d_inner * 2, bias=False)
+self.out_proj = $n$.Linear(self.d_inner, self.d_model, bias=False)
+self.gate_proj = $n$.Conv1d(self.d_model, self.d_inner, kernel_size=1, bias=True)
+# Components of F-SSM
+self.conv1d = $n$.Conv1d(self.d_inner, self.d_inner, kernel_size=d_conv,
+groups=self.d_inner, padding=d_conv - 1)
+self.x_proj = $n$.Linear(self.d_inner, self.dt_rank + self.d_state * 2, bias=False)
+self.dt_proj = $n$.Linear(self.dt_rank, self.d_inner, bias=True)
+self.A_log = $n$.Parameter(torch.log(torch.arange(1, d_state+1).repeat(d_inner, 1).float()))
+# Components of C-SSM
+self.conv1d_b = $n$.Conv1d(self.d_inner, self.d_inner, kernel_size=d_conv, groups=self.d_inner, padding=d_conv - 1)
+self.x_proj_b = $n$.Linear(self.d_inner, self.dt_rank + self.d_state*2, bias=False)
+self.dt_proj_b = $n$.Linear(self.dt_rank, self.d_inner, bias=True)
+self.A_b_log = $n$.Parameter(torch.log(torch.arange(1, d_state+1).repeat(d_inner, 1).float()))
+# Components of D-SSM
+self.conv1d_d = $n$.Conv1d(self.d_inner, self.d_inner, kernel_size=3, groups=self.d_inner, padding=1)
+self.x_proj_d = $n$.Linear(self.d_inner, self.dt_rank + self.d_state*2, bias=False)
+self.dt_proj_d = $n$.Linear(self.dt_rank, self.d_inner, bias=True)
+self.A_d = $n$.Parameter(torch.log(torch.linspace(0.5, 2.0, d_state).repeat(d_inner, 1)))
+# Deformable Scan
+self.point_deform_module = DeformableScanForPointCloud( in_channels=self.d_model, num_neighbors=self.num_neighbors, enable_dt=self.enable_dt,)
+self.linear = $n$.Conv1d(self.d_model, self.d_inner, kernel_size=1, bias=True)
+self.TPFF = TPFF(dim=self.d_inner, groups=32)
+self.act = $n$.GELU()
+def forward(self, hidden_states, inference_params=None, coords=None, base_idx=None):
+input: hidden_states [B, N, D], coords: [B, N, 3], base_idx: [B, N]
+output: out [B, N, D], new_coord: [B, N, 3], new_order: [B, N, 1]
+batch, seqlen, dim = hidden_states.shape
+# Input projection
+xz = rearrange(self.in_proj.weight @ rearrange(hidden_states, "b l d -> d (b l)"), "d (b l) -> b d l", l=seqlen,)
+if self.mamba_type == "v3":
+# 1. F-SSM
+A = -torch.exp(self.A_log.float())
+out_fwd = mamba_inner_fn_no_out_proj(xz, self.conv1d.weight, self.conv1d.bias, self.x_proj.weight, self.dt_proj.weight, A, None, None, self.D.float(), delta_bias=self.dt_proj.bias.float(), delta_softplus=True,)
+# 2. C-SSM
+A_b = -torch.exp(self.A_b_log.float())
+out_bwd = mamba_inner_fn_no_out_proj( xz.flip([-2]), self.conv1d_b.weight, self.conv1d_b.bias, self.x_proj_b.weight, self.dt_proj_b.weight, A_b, None, None, self.D_b.float(), delta_bias=self.dt_proj_b.bias.float(), delta_softplus=True,)
+# 3. D-SSM
+new_point_feat,new_coord ,new_order = self.point_deform_module(coords=coords, feats=hidden_states, base_idx=base_idx)
+new_point_feat = new_point_feat.permute(0, 2, 1)
+new_point_feat = self.linear(new_point_feat)
+x_deform_conv = self.act(self.conv1d_d(new_point_feat))
+x_dbl_d = rearrange(x_deform_conv, "b d l -> (b l) d")
+x_dbl_d = self.x_proj_d(x_dbl_d)
+dt_d, B_d, C_d = torch.split(x_dbl_d, [self.dt_rank, self.d_state, self.d_state], -1)
+dt_d = self.dt_proj_d.weight @ dt_d.t()
+B_d = rearrange(B_d, "(b l) dstate -> b dstate l", b=batch, l=seqlen)
+C_d = rearrange(C_d, "(b l) dstate -> b dstate l", b=batch, l=seqlen)
+dt_d = rearrange(dt_d, "d (b l) -> b d l", l=seqlen)
+out_def = selective_scan_fn( x_deform_conv, dt_d, -torch.exp(self.A_d.float()), B_d, C_d, torch.ones(self.d_inner), delta_softplus=True)
+# Fusion
+out = self.TPFF(out_fwd, out_bwd.flip([-2]), out_def)
+gate = torch.sigmoid(self.gate_proj(hidden_states.permute(0, 2, 1)))
+out = out * gate
+out = F.linear(rearrange(out, "b d l -> b l d"), self.out_proj.weight, self.out_proj.bias)
+return out, new_coord, new_order